### v2

# Reinforcing the Diffusion Chain of Lateral Thought with Diffusion Language Models

Abstract
We introduce the Diffusion Chain of Lateral Thought (DCoLT), a reasoning framework for diffusion language models. DCoLT treats each intermediate step in the reverse diffusion process as a latent "thinking" action and optimizes the entire reasoning trajectory to maximize the reward on the correctness of the final answer with outcome-based Reinforcement Learning (RL). Unlike traditional Chain-of-Thought (CoT) methods that follow a causal, linear thinking process, DCoLT allows bidirectional, non-linear reasoning with no strict rule on grammatical correctness amid its intermediate steps of thought. We implement DCoLT on two representative Diffusion Language Models (DLMs). First, we choose SEDD as a representative continuous-time discrete diffusion model, where its concrete score derives a probabilistic policy to maximize the RL reward over the entire sequence of intermediate diffusion steps. We further consider the discrete-time masked diffusion language model – LLaDA, and find that the order to predict and unmask tokens plays an essential role to optimize its RL action resulting from the ranking-based Unmasking Policy Module (UPM) defined by the Plackett-Luce model. Experiments on both math and code generation tasks show that using only public data and 16 H800 GPUs, DCoLT-reinforced DLMs outperform other DLMs trained by SFT or RL or even both. Notably, DCoLT-reinforced LLaDA boosts its reasoning accuracy by +9.8%, +5.7%, +11.4%, +19.5% on GSM8K, MATH, MBPP, and HumanEval.
1 Introduction
To enable complex reasoning, most large language models (LLMs) [15, 20] learn to decompose problems into simpler sub-steps and generate intermediate reasoning in natural language. Chain-of-Thought (CoT) [39] first reveals that step-by-step reasoning facilitates language models, as the outputs from previous steps could be rationales for more accurate next step prediction. Based on that, OpenAI’s PRM [23] supervises these intermediate reasoning steps with progressive rewards, to ensure the correctness of each single step. More recently, DeepSeek-R1 [15] eliminates the need of the reward model and verifies only the correctness of the final answer, relaxing the constraints on the reasoning process. However, due to the causal nature of attention mechanisms, auto-regressive models are still forced to reason in a single, sequential direction.
However, when developing ideas, human cognition does not always proceed through strictly sequential steps. At the beginning of thinking, human does not require an intact linguistic structure. Concepts, words, or ideas emerge spontaneously and independently first, and are gradually refined and organized over time to follow grammar rules. This non-linear and creative mode of reasoning, known as lateral thinking [17], contrasts with the structured, step-by-step approach of vertical thinking.
Figure 1: Comparison between CoT and DCoLT. (a) A typical CoT performs vertical thinking by following an auto-regressive convention that generates responses token by token from left to right in a linear way. (b) DCoLT performs lateral thinking that generates the responses in a non-linear way without following the auto-regressive order; moreover, at each step, it can generate multiple tokens at chosen positions. We focus on the lateral thinking in this paper by reinforcing the chain of such lateral thought as an entirety in Diffusion Language Models (DLMs).
Contrary to auto-regressive models, Diffusion Language Models (DLMs) [24, 27, 42] have also been adopted for text generation. The intermediate steps of the reverse diffusion process are naturally well-suited to emulate lateral thinking. Unlike auto-regressive models, diffusion models generate all tokens in parallel from a prior distribution. Each token can attend freely to all others under a non-causal mask in self attention, and intermediate reasoning steps are not required to conform to grammatical rules during multi-step generation, thus leading to more divergent thinking. In this paper, we propose the Diffusion Chain of Lateral Thought (DCoLT) to reinforce the lateral reasoning in diffusion language models, as shown in Fig. 1. Rather than providing explicit supervision for the thinking process, we employ outcome-based reinforcement learning, offering a rule-based reward that evaluates only the correctness of the final responses. This reward encourages the model to explore diverse, creative, and non-linear thought trajectories that ultimately lead to correct answers.
We study two paradigms of diffusion language models to reinforce the DCoLT, continuous-time diffusion language models [24, 6] and discrete-time diffusion language models [3, 19, 27, 42]. For the continuous-time paradigm, we consider SEDD [24] as a representative DLM. SEDD predicts the concrete score, allowing for a closed-form expression of the predicted diffused distribution of generated tokens at each step. This distribution can be viewed as a probabilistic policy for sampling tokens, which can be trained to optimize the reward on the final answers by reinforcing DCoLT.
For the discrete-time paradigm, we consider a masked DLM – LLaDA [27]. Besides the output distributions over discrete tokens, we note that the unmasking order plays an important role to decide which tokens ought to be kept to form the current step of lateral thought, and thus should be part of the learnable reasoning process. To this end, we introduce a Plackett–Luce model [32, 28] to define a ranking-based unmasking policy, where each masked token is assigned a predicted ranking score, and the unmasking policy selects the top-$K$ ranked tokens to retain in the output sequence at each diffusion step. The unmasking policy is trained together with the token generation policy to optimize the reward on the final answers.
We conduct experiments on both Math and code generation tasks to demonstrate the efficacy of DCoLT. After training the DCoLT on the SEDD 400M model, it achieves 96.2% and 57.0% in accuracy on Sudoku 4$\times$4 and GSM8K-Aug [10] tasks. On the LLaDA 8B model, DCoLT achieves the state-of-the-art performance among existing DLMs trained with SFT or RL or even both. Using only public data and 16 H800 GPUs, the DCoLT-reinforced LLaDA model achieves 88.1% on GSM8K [8], 44.6% on MATH [16], 51.6% on MBPP [4] and 59.1% on HumanEval [7] in the challenging zero-shot setting. Even compared with auto-regressive models that are trained with significantly more proprietary data and fully annotated CoT reasoning processes, it still demonstrates competitive performances.
2 Related Works
Diffusion Language Models
Diffusion models have achieved impressive results in image modeling [18, 37, 33, 21, 43], with recent efforts [24, 11, 13, 3, 48, 2, 47, 12, 29, 27, 42] extending to language tasks. Depending on the types of diffused distributions over text tokens, diffusion language models can be categorized into continuous diffusion models [13] and discrete diffusion models [3, 24, 11], with our work focusing primarily on the latter due to the discrete nature of language.
Among discrete diffusion models, masked diffusion models [27, 42, 24] emerge as a promising approach due to their superior performances. Recent advances simplify model training and design by adopting the straight cross-entropy loss [48, 29] and removing the time embeddings [47]. The resultant models [27, 42] achieve the state-of-the-art performance among existing diffusion language models.
Particularly, we consider two representative examples of discrete diffusion language models in this paper: SEDD [24] – a discrete diffusion model with a continuous diffusion time, and LLaDA [27] – a masked diffusion model with discrete diffusion timesteps. We will reinforce the entire sequence of their reverse diffusion processes containing non-linear text generations to release their lateral thinking ability. In contrast, DoT (Diffusion of Thought) [41] uses annotated step-by-step CoT data for supervised fine-tuning with existing diffusion losses such as the score entropy loss [24] or the noise prediction loss [13]. Thus, it still encourages a conventional stepwise reasoning process as in vertical thinking [17].
Reinforcement Learning for Language Models
Reinforcement Learning (RL) helps language models better align with human preference [30] or verifiable knowledge [22] (e.g. rewards on math/code tasks). Recently, there appear various algorithms for RL optimization [35, 1, 36, 25]. These methods enhance model’s ability to produce aligned outputs and show strong potential in inference-time scaling for challenging problems [15, 20].
However, existing RL approaches are primarily applied to auto-regressive language models [30, 15]. For DLMs, previous works explored policy gradient methods within the concrete score-matching framework [45], or estimated the log-probability of the sampled responses to perform reinforcement learning [46]. However, both models merely focused on the generated text response in the final diffusion step as an action to optimize, ignoring the role of intermediate diffusion steps as lateral thought to reinforce amid the reverse diffusion process.
Instead, in this paper, we seek to explicitly reinforce the reverse diffusion process to enable the training of lateral reasoning. We demonstrate in experiments that even though these intermediate diffusion steps perform non-linear rather than step-by-step reasoning, reinforcing them can eventually lead to correct answers at the final step that outperforms many CoT models [39, 10].
3 Methods
3.1 Diffusion Chain of Lateral Thought
To model the probability distribution $p_{\text{data}}$ over a finite vocabulary $\mathcal{V}=\{1,\dots,V\}$ for text generation tasks, discrete diffusion processes model how the unknown data distribution $p_{\text{data}}$ at $t=0$ gradually evolves into a prior distribution $p_{\text{prior}}$ at $t=T$ [6]. The distribution at an intermediate diffusion time $t$ is denoted by $p_{t}$.
Then a generation process is realized by reversing this diffusion process. Practically, it begins by sampling $x_{0}$ from the prior $p_{\text{prior}}$ and then iteratively denoising through a sequence of diffusion time $t_{0:N}$. At each step $n$, the model $\theta$ estimates the diffused distribution $p_{\theta,t_{n}}$ at time $t_{n}$, from which an intermediate sample $x_{n}$ is drawn, as shown in Eq. 1. The diffusion time $t_{n}$ decreases progressively as $n$ increases. After completing all denoising steps, reaching $t_{N}=0$, the final response $x_{N}$ is obtained.
$x_{{n}}\sim p_{\theta,t_{n}}(\cdot|x_{n-1})$ (1)
This naturally generates a series of intermediate results $x_{0:N-1}$ before arriving at the final output $x_{N}$. This behavior is analogous to the Chain-of-Thought (CoT) technique [39]. However, unlike the vertical and causal structure of CoT, the diffusion process enables the model to produce intermediate content that facilitates reaching final answers, aligning more closely with the concept of lateral thinking—solving problems through indirect and creative approaches. Thus, we define the entire sequence of all intermediate steps $x_{0:N}$ as the Diffusion Chain of Lateral Thought (DCoLT), as shown in Fig. 1, and seek to apply reinforcement learning to train it based on the given reward function.
Compared to standard CoT, DCoLT is distinguished with several notable features:
  • Bidirectional Reasoning: CoT generates tokens sequentially in a causal, irreversible manner; once derived, earlier thought remains fixed in the context. Instead, DCoLT adopts a bidirectional structure: each token is influenced by both preceding and succeeding content with bidirectional self-attention masks, allowing global refinement throughout generations.
  • Format-Free Reasoning: CoT typically adheres strictly to natural language format with complete linguistic structures. DCoLT, however, relaxes this constraint, allowing early stage of intermediate steps not necessarily to be complete or correct in format, thus enabling more divergent and creative patterns of thought before finally converging to complete responses.
  • Nonlinear Generation: CoT generates tokens one-by-one linearly in an auto-regressive manner from left to right. In contrast, DCoLT allows nonlinear generation of tokens at various positions. This aligns with how human develops ideas by beginning with keywords or critical points and then refining the details around them.
Fig. 7-Fig. 9 in Appendix B show some examples of intermediate responses from DCoLT-reinforced DLMs, demonstrating the above features with bidirectional and nonlinear generations throughout reverse diffusion processes from incomplete masked phases to complete unmasked responses.
To enhance the lateral thinking, we adopt final-outcome rewarded reinforcement learning to train the model, promoting reasoning trajectories that can lead to correct final responses $x_{N}$. For this, we treat the generation of $x_{1:N}$ as a sequence of actions to optimize as a whole. At each denoising step $n$, a distribution $\pi_{\theta,n}(\cdot|x_{n-1})$ over possible outputs is defined, which serves as the policy for sampling $x_{n}$ to calculate action probabilities during RL training. The reward $r$ is assigned based on verifiable correctness of final results. We do not impose any explicit supervision on the intermediate steps, promoting the model’s lateral thinking ability to explore diverse reasoning strategies to maximize the reward. This could yield useful patterns of thinking processes – for example, as shown in Fig. 3 of Appendix A, an easy-to-hard progressive generation of responses emerges from the DCoLT-trained DLM, in contrast to its counterpart supervisedly trained with the conventional diffusion loss.
We summarize the algorithm in Alg. 1. Details about training specific diffusion language models such as SEDD [24] and LLaDA [27] will be elaborated on in the following two subsections. Considering the demanding computation graph that expands multi-step generations in memory, we back-propagate the gradients at each step and accumulate them after the whole reverse diffusion process (c.f. Line 26 and 28 in Alg. 1). We use the GRPO [36] for fair comparisons with other RL-based models [36, 46], while alternative RL training approaches such as PPO [35] and RLOO [1] can also be adopted.
Algorithm 1 A General Framework for Training DCoLT
1:Model parameters $\theta$, a dataset $\mathcal{D}$, and reward_func.
2:while $\theta$ not converged and maximum epochs not reached do
3:    Sample questions $q\sim\mathcal{D}$
4:    for $g=1$ to $G$ do $\triangleright$ Generate a group of $G$ trajectories
5:         Initialize $x^{g}_{0}$ with $q$ and mask tokens.
6:         for $n=1$ to $N$ do $\triangleright$ $N$ denotes the number of denoising steps
7:             if training SEDD then
8:                 Sample $x^{g}_{{n}}\sim p_{\theta,t_{n}}(\cdot|x^{g}_{n-1})$
9:             else if training LLaDA then
10:                 Calculcate the ranking score $h_{\theta,n}$ for each token
11:                 Sample $K$ tokens to unmask in this step: $\mathcal{U}_{n}\sim\text{Plackett-Luce}(h_{\theta,n},K)$
12:                 Sample $x^{g,i}_{{n}}\sim p_{\theta,n}^{i}(\cdot|x^{g}_{n-1}),\qquad\forall i\in\mathcal{U}_{n}$
13:             end if
14:         end for
15:         $r^{g}=\operatorname{reward\_func}(q,x^{g}_{N})$triangleright Compute the rewards
16:    end for
17:    for $g=1$ to $G$ do $\triangleright$ Compute the advantages
18:         $A^{g}=\frac{r^{g}-\operatorname{mean}(r^{1:G})}{\operatorname{std}(r^{1:G})}$
19:    end for
20:    for $n=1$ to $N$ do $\triangleright$ Compute $\pi_{\theta}$ and losses for each denoising step
21:         if training SEDD then
22:             $\pi_{\theta,n}(x^{g}_{n}|x^{g}_{n-1})=\prod_{i=1}^{|x^{g}_{n}|}p_{\theta,t_{n}}(x_{n}^{g,i}|x^{g}_{{n-1}})$triangleright see Eq. 5
23:         else if training LLaDA then
24:             $\pi_{\theta,n}(x^{g}_{n}|x^{g}_{n-1})=\pi_{\theta,n}^{\text{unmask}}(\mathcal{U}^{g}_{n}|x_{n})\cdot\pi_{\theta,n}^{\text{token}}(x_{n}|x_{n-1},\mathcal{U}_{n})$triangleright see Eq. 9
25:         end if
26:         $\mathcal{L}_{\theta,n}=-\frac{1}{G}\sum_{g=1}^{G}\frac{\pi_{\theta,n}(x^{g}_{n}|x^{g}_{n-1})}{\pi_{\text{old},n}(x^{g}_{n}|x^{g}_{n-1})}A^{g}$
27:         Calculate the gradient $\nabla_{\theta}\mathcal{L}_{\theta,n}$
28:    end for
29:    Update $\theta$ with accumulated gradients $\sum_{n=1}^{N}\nabla_{\theta}\mathcal{L}_{\theta,n}$ along the descent direction
30:end while
3.2 A Continuous-Time DLM Case: DCoLT-reinforced SEDD
To define a discrete diffusion process, continuous-time diffusion language models such as SEDD [24] evolve a family of distributions $p_{t}$ according to a continuous-time Markov process, which can be represented by the following linear ordinary differential equation.
$\frac{dp_{t}}{d{t}}=Q_{t}p_{t},\quad p_{0}=p_{\text{data}},\quad p_{T}=p_{\text{prior}}$ (2)
Here, $Q_{t}$ is the transition rate matrix defining the forward process. We may first consider the simple single-token case ($Q_{t}\in\mathbb{R}^{V\times V}$). Conversely, to generate a sample from $p_{\text{prior}}$, this process has a corresponding reverse process, defined by a reverse transition rate matrix $\bar{Q}_{t}$:
$\frac{dp_{T-t}}{dt}=\bar{Q}_{T-t}p_{T-t},\quad\bar{Q}_{t}(y,x)=\frac{p_{t}(y)}{p_{t}(x)}Q_{t}(x,y),\quad\bar{Q}_{t}(x,x)=-\sum_{y\neq x}\bar{Q}_{t}(y,x)$ (3)
By Euler’s method, we have the transition probability $p_{t_{n}}$ at each diffusion step $t_{n}$, resulting in an iterative formula for multi-step generations from $t_{0}=T$ to $t_{N}=0$ with $n=0,\cdots,N$,
$p_{t_{n}}(x_{n}=y|x_{{n-1}}=x)=\delta_{xy}+(t_{n-1}-t_{n})\bar{Q}_{t_{n-1}}(y,x)$ (4)
with $x_{n}$ denoting $x_{t_{n}}$ to avoid notational clutter, i.e., the diffused sample $x_{t}$ at $t=t_{n}$.
Specifically, we consider a representative discrete diffusion model, SEDD [24]. This model learns to approximate the concrete score, i.e. $s_{\theta}(x,t)_{y}\approx\frac{p_{t}(y)}{p_{t}(x)}$ for any ${y\neq x}$ to represent the probability to transfer to other tokens. Thus, we may replace $\bar{Q}_{t_{n-1}}(y,x)$ with the model-estimated $s_{\theta}(x,t_{n-1})_{y}\cdot Q_{t_{n-1}}(x,y)$ in Eq. 4, making the transition probability tractable.
When scaling to sequences, we may apply $\tau$-leaping to define the probability of action at each timestep as the product of transition probabilities across all tokens:
$\pi_{\theta,n}(x_{n}|x_{n-1})$ $=\prod_{i=1}^{|x_{n}|}p_{\theta,t_{n}}(x_{n}^{i}|x_{n-1})$ (5)
$p_{\theta,t_{n}}(x_{n}^{i}|x_{n-1})$ =$=\begin{cases}s_{\theta}(x_{{n-1}},t_{n-1})_{i,x_{n}^{i}}\cdot(t_{n-1}-t_{n})\cdot Q_{t_{n-1}}(x_{n-1}^{i},x_{n}^{i}),&x_{n-1}^{i}\neq x_{n}^{i},\\ 1-\sum_{y\neq x_{n}^{i}}s_{\theta}(x_{{n-1}},t_{n-1})_{i,y}\cdot(t_{n-1}-t_{n})\cdot Q_{t_{n-1}}(x_{n-1}^{i},y),&x_{n-1}^{i}=x_{n}^{i},\end{cases}$ (6)
where $|x_{n}|$ denotes the token length of the sequence $x_{n}$. Note that the predicted $p_{t_{n}}$ is dependent on the model parameters $\theta$. During training, we may update $\theta$ to reinforce or suppress the probability of all actions along a trajectory of thoughts $[x_{0},x_{1},...,x_{N}]$, according to the reward associated with each completion. Unlike traditional diffusion model training, which typically optimizes each individual denoising step independently with ground truth responses, DCoLT jointly optimizes the entire reverse process of thought from $t_{0}=T$ to $t_{N}=0$, promoting the emergence of lateral thought leading to correct answers.
Figure 2: The structure of LLaDOU. It first predicts the token set to unmask $\mathcal{U}_{n}$ according to the ranking score $h_{\theta,n}$ by the UPM, and then samples those unmasked tokens in $\mathcal{U}_{n}$ by LLaDA blocks.
3.3 A Discrete-Time DLM Case: DCoLT-reinforced LLaDA
Some other diffusion language models instead operate at discrete timesteps [27, 42]. They can be viewed as Discrete-Time Diffusion Language Models with a multi-step generation process. For these models, we will figure out the probability of actions for each discrete step.
Among them is LLaDA [27] which is a typical Discrete-Time Masked Diffusion Language Model. It generates text by progressively unmasking a sequence initially prefilled entirely with mask tokens. At each generation step, the model receives a partially masked sequence as input. Then, it chooses and predicts a subset of masked tokens to reveal as clean text. As it progresses, the number of masked tokens gradually decreases, so that the model ultimately yields a fully generated sequence.
Specifically, we can define the model’s action at each step $n$ in two parts: 1) determining the set $\mathcal{U}_{n}$ of tokens to unmask, and 2) predicting the values of these tokens to obtain the new sequence $x_{n}$ over the unmasked part $\mathcal{U}_{n}$.
The action of selecting which tokens to unmask at each step can be decided by ranking masked tokens with a score function under the current state. To this end, we introduce an Unmask Policy Module (UPM), which predicts a score value $h_{\theta,n}^{i}$ for each token $i$ at the current diffusion step $n$. Based on these scores, we define a policy to sample a top-$K$ ranked list $\mathcal{U}_{n}=[u_{n}(1),...,u_{n}(K)]$ from a Plackett–Luce model [32, 28]: a multinomial distribution is formed from the predicted scores, and $K$ tokens are sequentially sampled without replacement, such that the corresponding scores are non-increasingly ordered $h_{\theta,n}^{u_{n}(1)}\geq\cdots\geq h_{\theta,n}^{u_{n}(K)}$ with high probability.
Formally, let $\mathcal{M}_{n}$ denote the set of tokens that remain masked after the $n$-th step, i.e., $\mathcal{M}_{n-1}\setminus\mathcal{U}_{n}$. Then, the probability of sampling a specific ranking list $\mathcal{U}_{n}$ is given by:
$\pi_{\theta,n}^{\text{unmask}}(\mathcal{U}_{n}|x_{n-1})=\prod_{k=1}^{K}\frac{\exp(h_{\theta,n}^{u_{n}(k)})}{\sum_{j=k}^{K}\exp(h_{\theta,n}^{u_{n}(j)})+\sum_{j\in\mathcal{M}_{n}}\exp(h_{\theta,n}^{u_{n}(j)})}$ (7)
Specifically, UPM takes the hidden states from the last hidden layer as inputs at each denoising step $n$, and predicts a ranking score $h_{\theta,n}^{i}$ for $i$-th token. It contains only one transformer block, introducing marginal computation cost. Moreover, the step index $n$ and the mask indicator of each token in $x_{n-1}$ can benefit the model training. Thus, we embed both of them into the Unmask Policy Module (UPM) via adaptive layer normalization. The module structure is shown in Fig. 2. For convenience, we name the DCoLT-trained LLaDA with UPM by LLaDOU (LLaDA with Ordered Unmasking).
As shown in Fig. 10 of Appendix B, the learned scores $h_{\theta,n}$ can be viewed as the ranked confidences over the token predictions at the current step $n$. Higher scores indicate that the UPM predicts lower levels of diffusion noises may be present in the generated tokens, which are less likely to contain errors and thus could be unmasked with higher priorities at the current step. Fig. 7 shows the resulting unmasking orders from the learned ranking scores for some prompts.
Once the unmask set $\mathcal{U}_{n}$ is determined, the model predicts their token values based on the output distribution by LLaDA blocks over the vocabulary. Viewing this prediction as a second-stage action, the probability of generating $x_{n}$ given $x_{n-1}$ and $\mathcal{U}_{n}$ can be written as
$\pi_{\theta,n}^{\text{token}}(x_{n}|x_{n-1},\mathcal{U}_{n})=\prod_{i\in\mathcal{U}_{n}}p_{\theta,n}(x_{n}^{i}|x_{n-1}).$ (8)
Thus, the probability of the complete policy for transitioning from $x_{n-1}$ to $x_{n}$ is the product of those of the unmask policy and the token prediction policy as used in Line 23 of Alg. 1:
$\pi_{\theta,n}(x_{n}|x_{n-1})=\pi_{\theta,n}^{\text{unmask}}(\mathcal{U}_{n}|x_{n})\cdot\pi_{\theta,n}^{\text{token}}(x_{n}|x_{n-1},\mathcal{U}_{n}).$ (9)
4 Experiments
We conduct DCoLT experiments on two DLMs – SEDD 400M and LLaDA 8B, each representing one of the two categories mentioned in Sec. 3.
4.1 Experiments on SEDD+DCoLT
4.1.1 Settings
We first conduct experiments on SEDD [24] for two different tasks: 1) Sudoku $4\times 4$ task and 2) GSM8K-Aug math problems [10], both assessing math reasoning performance. These two tasks allow us to perform direct comparisons with existing reasoning algorithms, including both supervised fine-tuning [10, 41] and reinforcement learning approaches [36]. For all experiments, we use the rule-based reward function to compute the reward for each sample: a reward of 1 is assigned if the solution is correct, and 0 otherwise. We choose SEDD as the base model to conduct DCoLT training, which is of medium size around 400 M. For detailed settings please refer to Appendix C.1.
4.1.2 Results
From the results in Tab. 1, our method remarkably outperforms both (implicit) CoT [10] and DoT [41] on both tasks. While both CoT and DoT perform vertical thinking, DoT also uses the SEDD as its base model to simulate causal reasoning. In contrast, our approach only uses RL to train the model without any CoT data for supervised training. For CoT, we adopt an auto-regressive language model – GPT2 [31] as the base model, which has the similar model size of about 400M to the SEDD model.
Particularly, SEED + DCoLT achieves an accuracy of 96.2% on Sudoku $4\times 4$ task, the best performance among all the methods. Using the same SEED model, DCoLT greatly exceeds DoT ($79.4\%$), even though the latter introduce step-by-step annotated CoT data for supervised training. For GPT2 model, both CoT and Implicit CoT, no matter if being post-trained with SFT or RL, perform worse than our model clearly. On the GSM8K-Aug dataset, SEED + DCoLT also reaches $57.0\%$ in accuracy, which is notably higher than both DoT and CoT counterparts. We will further analyze the generation process of SEED + DCoLT in Appendix A, demonstrating how tokens are generated in a lateral thinking manner.
Table 1: Experimental Results on SEDD. All models listed below are of medium size, with approximately 400M parameters. On the Sudoku $4\times 4$ dataset, we report the accuracy on a test set of 2,000 samples each with 9 empty cells. On GSM8K-Aug, we use the test set from GSM8K[8] for evaluation following [10].
Models Post-Training Sudoku 4$\times$4 GSM8K-Aug
GPT2 + CoT[10] SFT 71.5 43.9
GPT2 + Implicit CoT[10] SFT - 21.9
GPT2 + CoT RL 74.6 -
SEDD + DoT [41] SFT 79.4 53.5
SEDD + DCoLT RL 96.2 57.0
4.2 Experiments on LLaDOU (LLaDA+DCoLT)
4.2.1 Settings
We further apply DCoLT to LLaDA [27], a SOTA discrete-time masked-based DLM with 8B parameters. During generation, it starts with a masked sequence of length $256$, and performs a 256-step reverse diffusion process. We adopt a block-wise unmasking strategy as in LLaDA [27], and divide the sequence into blocks of length $8$. In Tab. 2, both LLaDA and LLaDOU (LLaDA+DCoLT) are evaluated with this setting for fair comparison. We conduct experiments with 16 H800 GPUs to jointly train both UPM and LLaDA parts. For implementation details, please refer to Appendix C.2.
4.2.2 Results
Table 2: Model accuracies on math and code generation benchmarks. The "Post-Training" column indicates what kinds of post-training (SFT or RL) phase the model goes through. indicates the model uses additional proprietary training data for the post-training phase. The numbers in parentheses represent the number of shots for the in-context learning, with “-” indicating unknown cases not mentioned in original papers. The results denoted with are evaluated with the prompt templates in Appendix C.2 for fair comparison, while others are reported in original papers. The results denoted with are evaluated on a subset MATH-500 instead of MATH. We highlight the best-performing model among compared DLMs in bold.
Method Post-Training Math Code
GSM8K MATH HumanEval MBPP
Diffusion Language Models
Dream 7B [42] baseline 81.1 (0) 42.9 (0) 51.8 (0) 49.6 (0)
LLaDA 8B [27] baseline 78.3 (0) 38.9 (0) 39.6 (0) 40.2 (0)
   + SFT [46]    + SFT 81.1 (0) 34.8 (0) - -
   + diffu-GRPO [46]    + RL 81.9 (0) 39.2 (0) - -
   d1-LLaDA (SFT + diffu-GRPO)    + SFT + RL 82.1 (0) 40.2 (0) - -
LLaDOU (LLaDA + DCoLT) 8B    + RL 88.1 (0) 44.6 (0) 59.1 (0) 51.6 (0)
Auto-regressive Models
LLaMA2 7B [38] baseline 14.6 (0) 2.5 (0) 12.8 (0) 20.8 (3)
   MetaMath 7B [44]    + SFT 66.5 (0) 19.8 (0) - -
   CodeLLaMA-Instruct 7B [34]    + SFT - - 34.8 (0) 44.4 (3)
Deepseek 7B [5] baseline 63.0 (0) 15.8 (0) 48.2 (0) 35.2 (3)
   DeepseekMath-Instruct 7B [36]    + SFT 82.9 (-) 46.8 (-) - -
   DeepseekMath-RL 7B [36]    + SFT + RL 88.2 (-) 51.7 (-) - -
   DeepseekCoder-Instruct 7B [14]    + SFT - - 66.1 (-) 65.4 (-)
As shown in Tab. 2, among all compared DLMs, LLaDOU consistently achieves the best performance across all benchmarks. On GSM8K, LLaDOU reaches 88.1%, significantly higher than other methods. On the more challenging MATH dataset, LLaDOU achieves 44.6%, outperforming baseline models such as LLaDA 8B (+5.7%) and Dream 7B (+1.7%), as well as models enhanced with post-training by SFT, RL or both. Notably, d1-LLaDA [46] uses 1K questions paired with detailed reasoning traces [26] in SFT to boost reasoning capabilities. In contrast, LLaDOU achieves superior performance – +6.0% on GSM8K and +4.4% on MATH – without any reasoning supervision, relying solely on reward signals based on the final answers.
Even when compared with auto-regressive models trained with a significantly larger amount of proprietary data, LLaDOU remains highly competitive. For example, DeepseekMath [36] gathers 776K questions with reasoning process annotations for SFT and 144K questions for RL, while LLaDOU is trained on just 15K public training samples from GSM8K and MATH. Despite using nearly two orders of magnitude fewer data for RL training, LLaDOU matches DeepseekMath’s performance on GSM8K (88.1%), highlighting its data efficiency and the effectiveness of reward-driven reasoning via diffusion-based lateral reasoning. In Appendix B, we analyze the thinking process generated by LLaDOU.
Similar trends are observed on code generation benchmarks. Unlike models that rely on valuable ground-truth code for supervised training, LLaDOU requires no code for supervised training at all. Instead, it only needs several test cases to provide a simple outcome-based reward for reinforcement learning: the model is rewarded only if the generated code passes all unit tests. Despite this code-free reward signal, LLaDOU achieves 51.6% on MBPP and 59.1% on HumanEval, outperforming other DLMs and auto-regressive models, except for DeepseekCoder [14] that benefits from 2B tokens of high-quality instruction-tuning data—approximately 150×150$\times$150 × more than we used (about 12M tokens). Details of post-training datasets used by all compared methods are provided in Appendix C.3.
4.2.3 Ablation Studies
We provide some ablation studies to reveal the role of some model design and hyper-parameter settings in LLaDOU. Unless otherwise specified, all experiments are conducted on LLaDOU 8B with $N=64$ diffusion steps, and the model is trained for $150$ iterations with a batch size of $32$.
Better performance with the UPM
Table 3: Ablation on the Unmasking Policy Module
Trained parameters GSM8K Acc.
UPM LLaDA
$\times$ $\times$ 47.27 (Baseline)
w/ AdaLN $\times$ 69.24
w/o AdaLN 80.53
w/ AdaLN 81.06
In Tab. 3, we ablate the model design in LLaDOU. Even if we freeze the original model parameters in LLaDA part and only train the UPM part, it achieves a significant improvement from 47.27% to 69.24%. This result indicates that the unmasking policy plays a crucial role in our model. By training the LLaDA part with RL together, the accuracy further improves to 81.06%.
We use the adaptive layernorm in the UPM. Removing it degrades accuracy to 80.53%. This demonstrates that incorporating the embeddings of diffusion step $n$ and the mask indicators also benefits the training of the unmasking policy.
Extension to longer generation length
Though the models in Tab. 2 is trained to generate completions with a fixed generation length $256$ (i.e., the length of initialized mask tokens), it can benefit from generating longer sequences without further RL training on longer generations. This is especially effective when answering difficult questions on MATH dataset, in which solving problems sometimes require longer reasoning. Tab. 4 reports the accuracies with different generation lengths. It suggests that longer generations in inference lead to better accuracies. For example, a generation length of $384$ achieves an accuracy of 46.22%, which is +1.62% higher than the default length of $256$. This improvement vanishes as we further increase the generation length. However, if we further tune the model on a longer generation length, for example $L=512$, the accuracy further improves to 47.3%. This suggests the potential of how scaled generation lengths may improve the performance of DCoLT-trained DLMs, as shown in Fig. 13 of Appendix D.
Table 4: Model accuracies with extended generation lengths on MATH without further RL-finetuning the model on these longer generations. The results denoted with are taken from [46] and evaluated on a subset MATH-500 instead of MATH.
Model Generation Length
256 320 384 448 512
LLaDA 8B [27] 38.9 40.1 41.5 42.3 42.5
   + SFT [46] 38.6 - - - 40.2
   + diffu-GRPO [46] 37.2 - - - 39.2
   d1-LLaDA (SFT+diffu-GRPO) [46] 38.6 - - - 40.2
LLaDA + DCoLT (LLaDOU) 8B 44.6 45.7 46.2 45.7 45.9
5 Conclusion
In this paper, we propose a new paradigm of model reasoning ability, Diffusion Chain of Lateral Thought (DCoLT), for diffusion language models. DCoLT considers the intermediate steps in the reverse diffusion process as the thinking actions, and optimizes this process with final-outcome rewarded Reinforcement Learning (RL). We implement DCoLT on two representative diffusion language models, SEDD and LLaDA. On SEDD, we derive the corresponding RL optimization over the distribution chain of thinking actions via the predicted concrete scores at denoising steps. On the masked-based diffusion language model LLaDA, we regard the order in which tokens are unmasked at each diffusion step as its action, and propose an Unmask Policy Module (UPM) to optimize over the order distribution with the Plackett-Luce model. The experiments indicate that DCoLT outperforms other SFT- or RL-based training algorithms, and demonstrate its effectiveness on a wide range of tasks and benchmarks, including math and code generation.
Limitations
First, due to limited training data and compute, our model’s performances on Math and code generation tasks still have much rooms to improve. Prior works [36, 44, 14, 34] demonstrate that proprietary data often significantly improves the model performances. Also, scaling the token length of sequences during training can also enhance reasoning. We will seek to boost the model ability along these directions when computing resources become available. Second, currently DCoLT is only validated on tasks having a verifiable reward function. We may need a reward model to cover more general tasks. We will develop these further in our future research.
Appendix A Analysis of Thinking Process in SEDD+DCoLT
For the Sudoku dataset, we aim to visualize the thinking process by analyzing the generation step of different cells. We categorize all $4\times 4$ cells into three groups: (1) given cells – those initially provided by the puzzle; (2) easy cells – those that can be directly inferred using basic Sudoku rules (each number 1–4 must appear exactly once per row, column, and $2\times 2$ subgrid); and (3) hard cells – the remaining positions that require more complex reasoning, often involving the resolution of dependencies among other non-given cells.
In Fig. 3, we plot the distribution of generation step for these three cell types, the left side displays statistics for the SEDD + SFT model, while the right side shows our further RL-trained SEDD + DCoLT model. The right figure reveals that the our SEDD + DCoLT significantly prioritizes generating easy cells in earlier steps than hard cells, as these cells can be directly determined from the initial given cells, indicating that our model learns a progressive generation strategy from easy to hard. Notably, in the left figure, we observe almost no difference between generation orders across various difficulty levels for SEDD + SFT model. This comparison reveals that the DCoLT changes its SFT-trained counterpart’s reasoning behavior to more progressive generation aligned naturally with how humankind handles problems in a real world.
Figure 3: This figure shows the model’s inference patterns on Sudoku dataset, with the left side displaying the pattern of SEDD + SFT model and the right side showing that of our SEDD + DCoLT model. We plot the total number of generated tokens for these three cell types over diffusion steps on Sudoku $4\times 4$ dataset, with the number of sampling steps set to 32.
To further demonstrates the progressive generation behavior on sample level, we visualize the generation process of our SEDD + DCoLT on the Sudoku dataset in Fig. 4. We show the predicted $\hat{x}_{0}$ at each timestep, obtained by selecting the token with the highest probability (excluding the mask token) at each token position. As shown in the top row, positions (0, 0) and (0, 1) can be regarded as hard cells, because the value can’t be directly inferred from given cells denoted by black borders. Initially, our model assigns the highest probabilities to incorrect answers in these red cells due to insufficient contextual information. However, as the denoising process progresses, the model gradually refines its predictions by leveraging information from newly unmasked cells (shown in dark blue), eventually converging to the correct values. This demonstrates the importance of nonlinear generation for Sudoku $4\times 4$ tasks.
Figure 4: Visualization of predicted tokens $\hat{x}_{0}$ by SEDD + DCoLT on Sudoku $4\times 4$: those that are still masked appear in light blue, with unmasked ones in dark blue, incorrect predictions in red, and corrected ones in green. Black borders indicate the given cells.
For GSM8K-Aug dataset, since it’s hard to determine which token position is more difficult to predict, we only show the generation process and the generation step statistics for each position. Fig. 5 demonstrates the contrasting generation behaviors between SEDD + DCoLT and GPT2 + CoT on the GSM8K-Aug dataset. While GPT2 + CoT consistently follows a left-to-right generation order (where earlier generation steps correspond to lower token positions), SEDD + DCoLT shows minimal positional bias in generation order across token positions. Furthermore, our results in Fig. 6 reveal that SEDD + DCoLT’s generation order is sample-dependent, maintaining a nonlinear generation throughout the process.
Figure 5: Average generation step for each token position in GSM8K-Aug. Different from CoT, SEDD + DCoLT generates in a non-linear way.
Figure 6: Our model’s generation process on GSM8K-Aug, each line displaying the predicted-$\hat{x}_{0}$ tokens across different steps, with masked tokens shown in gray and unmasked ones shown in black.
Appendix B Analysis of Thinking Process in LLaDOU (LLaDA+DCoLT)
We illustrate how LLaDOU generates a 64-token answer to a GSM8K question in Fig. 7. Tokens shown in darker shades are generated in later diffusion steps. As the figure shows, key numbers and symbols tend to emerge early in the generation process, while surrounding textual elements are filled in later to ensure grammatical correctness and fluency.
Fig. 8 and Fig. 9 further show intermediate steps of the output when generating responses of length 256 with 64 steps, in which key numbers and symbols are generated to gradually form the structure of thinking processes before grammatically correct sentences are completed.
Fig. 10 visualizes the predicted ranking scores $h_{\theta,n}$ to unmask tokens during the generation process. The examples show that at each step, the unmasking score evaluates how likely the predicted token is correct at each position. For example, in Step 20 of case (a), the token “3" prior to the selected token “0" has a smaller value of the unmasking score, suggesting that it is more likely than “0" to be wrong. This token is corrected to “0" before being unmasked at a later step.
(a) Miguel uses 2 pads of paper a week for his drawing. If there are 30 sheets of paper on a pad of paper, how many sheets of paper does he use every month?
(b) Two trains leave San Rafael at the same time. They begin traveling westward, both traveling for 80 miles. The next day, they travel northwards, covering 150 miles. What’s the distance covered by each train in the two days?
(c) Janet buys a brooch for her daughter. She pays \$500 for the material to make it and then another \$800 for the jeweler to construct it. After that, she pays 10% of that to get it insured. How much did she pay?
(d) An 8-year old child wants to buy a toy car which costs \$12. He already has \$4 savings. How many days will it take him to save the remaining amount of money if he promises to save \$2 daily from his allowance?
Figure 7: Examples of 64-token responses generated by LLaDOU on GSM8K. Token color reflects the generation order, progressing from light (early steps) to dark (later steps). The corresponding questions are provided in the sub-captions. We observe from the results that many tokens bearing key information to final answers such as the numbers, units and factual entities tend to be unmasked at earlier steps, while those tokens for function words and symbols such as “of" and “the" tend to be unmasked at later steps.
Figure 8: Intermediate step outputs from LLaDOU during the generation of a 256-token response. The question is from GSM8K: A robe takes 2 bolts of blue fiber and half that much white fiber. How many bolts in total does it take?
Figure 9: Intermediate step outputs from LLaDOU during the generation of a 256-token response. The question is from GSM8K: Cars have lined up on the motorway. Some of the cars drive through in the first 15 minutes of the traffic jam, then 20 more cars drive through in the remaining 15 minutes of the jam. 5 cars from the line take an exit so they don’t have to drive through the traffic jam. If there were originally 30 cars on the motorway, how many cars drove through the traffic jam in the first 15 minutes?
(a) The number of students in a school hall was 1000. The hall had 3 entrances A, B, and C which also served as the exits. If after a meeting 30% of the students went out of the building through exit A, 3/5 of the remaining went out through exit B, and the rest went out through exit C, calculate the number of students who went out through exit C.
(b) James has to buy insurance. Since he had an accident it was 60% more than normal. The normal cost is \$120 a month. How much does he pay a year?
Figure 10: Illustration of intermediate steps of generated responses by LLaDOU on GSM8K. Gray tokens represent positions that have already been unmasked prior to the current step. For each masked position, we show the predicted token with the highest generation probability. Token color indicates the unmasking score, with darker blue indicating higher probability of be unmasked. The token eventually selected to unmask at each step is bounded in a red box. The corresponding prompts are given in the sub-captions.
Appendix C Experiment Settings
C.1 SEDD+DCoLT
Table 5: Evaluation Hyperparameters for Sudoku and GSM8K-Aug
Models Sudoku GSM8K-Aug
Generation Tokens NFEs Temp. Generation Tokens NFEs Temp.
GPT2 + CoT 256 256 0.0 256 256 -
SEDD + DoT 256 512 0.5 256 64 0.5
SEDD + DCoLT 19 32 0.5 64 64 0.0
Table 6: Training Hyperparameters for Sudoku
Models Post-Training Post-Training Steps Learning Rate Batch Size
GPT2 + CoT SFT 1,500 3e-4 1024
SEDD + DoT SFT 1,500 1e-4 1024
GPT2 + CoT RL 500 1e-5 64
SEDD + DCoLT RL 500 1e-5 64
Sudoku $4\times 4$
We first generated 50,000 sudoku puzzles of size $4\times 4$ by uniformly replacing 1-9 cells with zeros. Using DeepSeek V3-0324[9], we produced CoT reasoning steps and filtered the results to remove repetitive responses and incorrect solutions. All training datasets were created from this cleaned dataset to ensure fair comparison. Dataset samples can be found in the Fig. 11.
Figure 11: Examples of CoT data used for training GPT2+CoT and SEDD+DoT.
We trained both DoT and CoT (with SFT) for 1,500 steps. For CoT (with RL), we first performed 1,000 steps of SFT to pretrain the CoT response, reaching 70.5% in accuracy, followed by another 500 steps of RL training using GRPO[36]. For a fair comparison, DCoLT was also pre-trained to reach 72.3% accuracy before applying 500 steps of RL training. Tab. 6 presents training hyperparameters for the Sudoku $4\times 4$ dataset. For both SFT and RL post-training, we employ Adam optimizer with $(\beta_{1},\beta_{2})=(0.9,0.999)$. During RL post-training, we set the group size of GRPO to 32, and the kl coefficient to 0.
GSM8K-Aug
Our experiments are conducted on the GSM8K-Aug dataset (sourced from [10]), which contains 384,623 training samples with simplified CoT processes. After filtering out samples with too long answers (exceeding 64 tokens) and those with invalid formats, we retained 382,553 examples. For the pretraining of SEDD on GSM8K-Aug, we train it for 120K steps using the Adam optimizer with a learning rate of $3\times 10^{-4}$, a batch size of 512, and a generation length of 64 tokens. In the RL training phase, we employ the Adam optimizer with a learning rate of $5\times 10^{-5}$ and a batch size of 32, and set the group size for GRPO to 64. This RL phase is trained for 1,000 steps.
All training was conducted on 8 H800 GPUs. For GSM8K-Aug pre-training, we spent 160 GPU hours on pretraining for 120K steps, and additional 168 GPU hours for the RL post-training. Evaluation hyperparameters for both datasets are provided in Tab. 5.
C.2 LLaDOU (LLaDA+DCoLT)
C.2.1 Implementation Details for Math Problems
Datasets & Benchmarks
We consider two popular benchmarks, GSM8K [8] and MATH [16]. In our experiments, we follow the common train-test split on these datasets. For GSM8K, there are 7.5K questions for training and 1.32K questions for testing. For MATH, there are 7.5K questions for training and 5K questions for testing. We report the accuracy on their test set. During training, we extract the final answers from generated responses to compute the rewards.
Reward function
We assign a hard reward for each completion, that it is $1$ only if the final answer is equivalent to the ground truth [16].
Training configurations
The model is trained with 64 prompts in a batch, each generating 16 completions to form a group for advantage calculation. We take an AdamW optimizer with a learning rate of $5\times 10^{-6}$, and $(\beta_{1},\beta_{2})=(0.9,0.999)$. We do not apply the KL penalty by default, as it provides marginal benefits in our experiments. The whole training lasts for $140$ iterations on 16 H800 GPUs, which takes about 63 GPU days (i.e., about 4 days on wall clock with 16 GPUs).
C.2.2 Implementation Details for Code Generation
Dataset
We filter the KodCode-V1-SFT-R1 [40] dataset to construct our training dataset. KodCode is a large fully-synthetic open-source dataset providing verifiable code solutions and test cases for coding tasks.
To ensure the quality and consistency of the training data, we retain only samples labeled with the "instruct" style and containing exactly one entry function for test with a non-empty docstring. We further exclude any samples where the provided solutions include class definitions or constructor methods, and ensure that the solutions contain exactly one function definition. Additionally, we only use samples from a group of subsets, including Prefill, Taco, Leetcode, Codeforces, Code Contests, and Filter, for their appropriate difficulty level.
After filtering, we obtain a training set with 48.9K samples (around 12M tokens). We then format the samples into the prompt template below:
You are an expert Python programmer. Your task is to complete the implementation of a function named `<function_name>`.
** TARGET FUNCTION **
<docstring description>
** UNIT TESTS **
Your code should pass unit tests like:
      <assert statement 1>
      <assert statement 2>
Here is the function to complete:
```python
      def <function_name>(<parameters>):
            """<docstring description>"""
Reward function
We assign a hard 0/1 reward for each completion, that it is $1$ only if the generated code block passes all provided test cases. The code blocks are extracted from the responses based on the markdown format.
Training configurations
The training configurations are kept the same as for math. The training lasts for 240 iterations, and it takes about 127 GPU days on 16 H800 GPUs.
Evaluation details
We evaluate the model in a zero-shot setting on MBPP [4] and HumanEval [7], and report the pass@1 metric. We show some examples of their input prompts below.
  • MBPP:
    You are an expert Python programmer.Your task is to complete the implementation of a function named `remove_Occ`.
    ** TARGET FUNCTION **
    Write a python function to remove first and last occurrence of a given character from the string.
    ** UNIT TESTS **
    Your code should pass unit tests like:
          assert remove_Occ("hello", "l") == "heo"
          assert remove_Occ("abcda", "a") == "bcd"
          assert remove_Occ("PHP", "P") == "H"
    Here is the function to complete:
    ```python
    def remove_Occ(input_param_1, input_param_2):
          """Write a python function to remove first and last occurrence of a
          given character from the string."""
  • Humaneval:
    You are an expert Python programmer.Your task is to complete the implementation of a function named `has_close_elements`.
    Here is the function to complete:
    ```python
    from typing import List
    def has_close_elements(numbers: List[float], threshold: float) -> bool:
          """Check if in given list of numbers, any two numbers are closer to each
          other than the given threshold.
          Examples:
          $>$ has_close_elements([1.0, 2.0, 3.0], 0.5)
          False
          $>$ has_close_elements([1.0, 2.8, 3.0, 4.0, 5.0, 2.0], 0.3)
          True
C.3 Post-Training Data Used in Tab. 2
In Tab. 7, we list the data used in post-training for math and code generation tasks by different models. Note that some methods [34, 14, 36] involve a considerable large amount of proprietary math/code data for both pretraining and post-training of models, which makes crucial contributions to high accuracies.
Table 7: Post-training data for math / code generation used in Tab. 2. “-" means this type of post-training was not used for the model.
Method SFT RL
Auto-regressive Models
LLaMA2 7B [38] - -
   MetaMath 7B [44] 395K samples (MetaMathQA) -
   CodeLLaMA-Instruct 7B [34] thousands of SFT + millions of RS examples -
Deepseek 7B [5] - -
   DeepseekMath-Instruct 7B [36] 776K (problem, solution) pairs -
   DeepseekMath-RL 7B [36] 776K (problem, solution) pairs 144K questions
   DeepseekCoder-Instruct 7B [14] 2B tokens -
Diffusion Language Models
Dream 7B [42] - -
LLaDA 8B [27] - -
   + SFT [46] 1K samples from s1K -
   + diffu-GRPO [46] - GSM8K: 7.5K / MATH: 7.5K
   d1-LLaDA (SFT + diffu-GRPO) 1K samples from s1K GSM8K: 7.5K / MATH: 7.5K
LLaDOU (LLaDA + DCoLT) 8B - Math: 7.5K GSM8K + 7.5K MATH / Code: 48K filtered from KodCode
Appendix D Potential Scaling Law for Longer Generations
Fig. 13 illustrates some examples of generated responses of various lengths to the same prompt for the LLaDOU. In DLMs, the generation length tends to increase as more mask tokens are initialized at the beginning of the reverse diffusion process. This differs from auto-regressive language models, where the length of a generated response is determined by when the End of Text (EoT) token appears.
Although DLMs also use the EoT token to mark the end of the output, the token often appears later in the output sequence to generate a longer thinking process when more mask tokens are initialized. This offers a flexible way for DLMs to control various lengths of generated responses. As shown in Tab. 4, longer generations could improve performance, particularly when the model is fine-tuned by DCoLT with increasing generation length. This suggests the existence of a potential scaling law for longer generations.
Fig. 13(a) shows that some mistakes generated in the thinking process can be corrected in longer generations, thus achieving better results over shorter ones. We will study such a scaling phenomenon further in the future.
(a) The graphs of $y=x^{4}$ and $y=7x^{2}-10$ intersect at four points with $x$-coordinates ±mplus-or-$\pm\sqrt{m}$ and ±$\pm\sqrt{n}$, where $m>n$. What is $m-n$?
(a) Find $4321_{5}-1234_{5}$. Express your answer in base $5$.
Figure 13: Examples of how increasing generation lengths enable the DCoLT-reinforced model to unfold more complex reasoning processes. Incorrect steps are marked in red, while the corrected steps are highlighted in green. The generation length means the number of mask tokens initialized in the input sequence to the reverse diffusion process. Except for the model finetuned with 512 generation length, other models are only trained with 256 generation length as shown in experiments.
Appendix E Ablation on Various Block Lengths
Table 8: Model accuracies with different block length on GSM8K dataset. LLaDOU (LLaDA + DCoLT) is trained with a generation length of 256 and 64 steps of reverse diffusion under respective block lengths.
Model Block Length
8 32 256
LLaDA 8B 63.26 62.12 49.47
LLaDOU (LLaDA+DCoLT) 8B 83.17 83.11 82.27
LLaDA [27] divides the sequence into several blocks and generates responses blockwise from left to right in a semi-autoregressive manner. This policy introduces additional priors on the block-wise order of text generation. In Tab. 8, for the baseline LLaDA 8B model, setting block length to $8$ achieves an accuracy of 63.26% on GSM8K, significantly higher than 49.47% without dividing blocks.
However, when we apply DCoLT to LLaDA with different block lengths, all LLaDOU models achieve similar performance, with the one with a block length 8 achieving slightly higher accurcy (83.17%) than the one with block length 256 without dividing blocks (82.27%). These results suggest that LLaDOU does not rely on the blocking prior as in the baseline model.
Appendix F Licenses for existing assets
The code and models associated with this paper will be released to the public later. For all code and data assets, we refer to their licenses in Tab. 9.
Table 9: Reference assets and their licenses.
Asset License Utility
SEDD [24] MIT Code & Model
GSM8K-Aug [10] - Data
LLaDA [27] MIT Code & Model
MATH [16] MIT Data
GSM8K [8] MIT Data
KodCode [40] CC BY-NC 4.0 Data
Appendix G Broader Impacts
Our work introduces a novel perspective of lateral reasoning in language models, in contrast to the widely adopted CoT paradigms used in LLMs. By enabling non-linear, format-free reasoning through a reverse diffusion process, our approach offers new insights into the mechanisms of reasoning in generative models. This perspective has the potential to benefit a broad range of reasoning-intensive tasks, including embodied AI, autonomous agents, and complex decision-making systems, providing transparent thinking processes to minimize the hallucination risks that could emerge from inadequate reasoning of language models.
However, we acknowledge that DCoLT could pose potential unexpected societal impacts if misused, especially when handling misleading or adversarial prompts. In this work, we focus on math and code generation tasks, in which outputs are objectively verifiable. With such a reward design, we can reduce the influence of dataset biases and encourage the development of reasoning behaviors that are aligned with the factual and logical consistency. Handling subjective preferences in rewarding the models to train the DCoLT could be more challenging, and we leave it to our future works.
Appendix H Safeguards
DCoLT is designed for math and code generation tasks that involve objectively verifiable outputs and well-defined correctness criteria. These domains present a relatively low risk of societal misuse compared to open-ended language generation tasks. To further reduce the potential for unintended use, we recommend deploying DCoLT alongside an input checker that ensures that the model only processes input relevant to its target domains. This approach helps mitigate the risks related to adversarial prompting or misuse beyond the intended scope.